How to convert aniline to methylamine?
1
Expert's answer
2013-02-25T11:49:13-0500
Methylamine synthesis can be carried out using the following conversion:

At the first stage the methylation of aniline on the nitrogen atom:

Ph-NH2 + CH3I = Ph-NH-CH3 + HI with K2CO3

CH3I dropwise within one hour in an ice bath.

After the dropwise addition, the mixture is stirred at room temperature for about 24 hours.
The second stage - the nucleophilic substitution of a good leaving group strong nucleophile:

Ph-NH-CH3 + CH3OH = Ph-O-CH3 + CH3-NH2